OpenStudy (anonymous):

Find the dimensions of a rectangle whose perimeter is 46 m and whose area is 126 m^2.

OpenStudy (anonymous):

(L+w)2=46 L+w =46\2 L+W=23 L*W=23 14+9=13 14*9=126 14 and 9